# Are Triggers Needed for Document-Level Event Extraction?

### v1

@@ 5 Models
We study the performance of four state-of-the-art sequence-to-sequence style systems for document-level event extraction: TANL Paolini et al. (2021), GTT Du et al. (2021), Degree Hsu et al. (2022), and GenIE Li et al. (2021) and compare them against prompt based baselines using GPT-4o and GPT-4o-mini.
@@ 5.1 Models Overview
We prompt GPT-4o and GPT-4o-mini to extract event templates in a JSON format prompting it with six in-context examples. We extract all events for all event types using a single query. We use event descriptions and templates as instructions. An example of the prompt can be found in Appendix B.
@@
Paolini et al. (2021) is a pipeline-based model that takes only the document as input. As shown in Figure 2, TANL fine-tunes a T5-large Raffel et al. (2020) to complete event extraction in two phases : 1) event detection, and 2) argument extraction. First, the document is passed through T5 to tag/delimit a trigger for each detected event (the entire document is produced as output with event type tags associated with recognized triggers). Then, each identified {trigger, event type} pair (tagged/delimited in the input document) is passed through T5 once to identify all event arguments and assign them a role type.
@@
Du et al. (2021) is an end-to-end model that takes a document and an ordered list of the possible event types and generates templates for all event types in a single pass through BERT Devlin et al. (2019). Templates are linearized to sequences with special argument and event delimiters.
@@
Hsu et al. (2022) is an end-to-end model that takes in a document, a natural language template for event type E, extra information on events of type E (i.e. type definitions, example triggers) and generates filled templates (that include identified triggers) for all detected events for type E in a single pass. The model is called once for each event type of interest for each document. For this reason, training is augmented with multiple negative examples (i.e. document-(empty) template pairs for event types that do not appear in the text).
@@
Li et al. (2021) is originally proposed for argument extraction. Specifically, a document is provided with the trigger delimited by special tokens and a natural language template (similar to that of Degree) corresponding to the event type to extract arguments for. The model output is the template filled with arguments. We use GenIE for the full event extraction task by using TANL as the trigger extractor and GenIE as the argument extractor.
@@
Figure 3: E2E vs. Pipeline. This plots the F1-score of each e2e experiment against its pipeline version. If the data point is above $y=x$ (dashed line), then the e2e architecture outperformed the pipeline architecture for that given (model, trigger type, dataset) setting.
@@
Figure 4: Triggers vs. no-Triggers. This plots the F1-score of each fine-tuning experiment with triggers (annotated by humans, LLMs, or the keyword heuristic) against it’s e2e version without triggers. A data point above $y=x$ (dashed line) implies that fine-tuning with triggers performed better for that (model, architecture, dataset).
5.2 Model Type: Pipeline vs End-to-End (e2e), No Trigger
We explore both pipeline and end-to-end model architectures and modify TANL and Degree to run in both settings. As shown in Figure 2, a pipeline system consists of a trigger extraction and an argument extraction component. The trigger extraction phase identifies all event triggers and their corresponding event types. The argument extraction phase then operates on each {trigger, event type} pair separately. This design allows argument extraction to anchor in an informative part of the text, i.e. in the vicinity of a known trigger. End-to-End systems, on the other hand, jointly extract triggers, event types, and arguments, allowing the two extraction tasks to inform one another, thus alleviating cascading errors that can occur in a pipelined setting. All no trigger experiments were conducted in an e2e manner.
Pipeline
We modify Degree to fit the pipeline architecture. We train distinct models for trigger and argument extraction. Specifically, the trigger extraction input and output templates only contain slots for a trigger and event type; the argument extraction input templates have an additional slot for a trigger (identified during trigger extraction) and the output templates only remove the slot for the trigger, retaining only slots for each argument type. The model is called once to generate templates for all identified triggers (trigger extraction) and once for each {identified trigger, event type} pair to identify arguments (argument extraction). We did not use GTT for pipelined experiments. By default, TANL, TANL + GenIE are pipelined models.
No Triggers
We modify all models to run without the use of triggers. For TANL, the raw text is provided and the model tags arguments with an event type and number and the argument role (i.e. “perpetrating individual for attack 1”). A single span can have multiple such tags if it is a shared entity mention for multiple events. The spans associated with each unique {event type, number} pair are processed into fillers for single events. This scheme only requires one pass of T5 for full event extraction. For Degree, we only remove the slot for the trigger in input/output templates. By default, GTT is an e2e system that does not use triggers.
End-to-End (with triggers)
For TANL, the event type and number and the argument role tag is supplemented with a tag for the trigger (i.e. “perpetrating individual for attack 1 | trigger = killed”). There is still only one pass for full event extraction. For GTT, a slot is added at the start the output event templates for a trigger. By default, Degree is an e2e system using triggers.
6 Experimental Setup
In our study of the utility of triggers on document-level event extraction, we tested the performance of fine-tuned and prompt-based approaches on the three datasets described in Section 4. We consider the prompt-based approaches as a baseline and compare them to fine-tuning the four state-of-the-art models, described in Section 5; each model was tested in a pipeline and an end-to-end manner as applicable. We train on 1 $\times$ A100 and 4 $\times$ 1080Ti using the default learning rate suggested for each model but unified the warm-up for the first 10% of the steps. As we have different base models, we report the score on the test data from the best performing epoch against the validation set.
We evaluate all the experiments using overall Micro-F1 — combined F1 score on both event type detection and argument role extraction. Since the datasets are multi-event, we have to align the predicted and gold templates. We follow the Das et al. (2022) matching algorithm to align and score the predicted templates. We only modify their proposed scorer by improving its compute time (using the Hungarian Matching algorithm Kuhn (1955)) and by removing span-based scoring as not all models predict spans.
7 Results and Discussion
In this section, we discuss our experimental results shown in Table 2. We first analyze some high level performance trends independent of trigger usage before commenting on how triggers should be used depending on various system and corpora specifications.
7.1 General Observations for Document-Level Event Extraction
Both prompt-based LLM methods consistently fail to outperform fine-tuning of smaller-scale LLMs across all datasets. As shown in Table 2, the best prompt-based systems achieve 43.3%, 36.2%, and 58.0% F1 while the best fine-tuning methods achieve 56.3%, 43.3%, and 71.8% F1 on MUC, WikiEvents and CMNEE, respectively.
Across all datasets, all e2e models trained with triggers perform very comparably to their pipelined counterparts with only one exception. In Figure 3, we plot the F1-scores for each e2e model vs. its pipelined variant; that is, the same model trained on the same trigger source on the same dataset. We see that almost all points are near, or above, the $y=x$ dashed line, which shows that e2e models are comparable to, and sometimes better than, pipelined models. This is important because we can generally build e2e systems that are faster to train and quicker during inference than pipelined systems without sacrificing performance. The exception to this is the TANL system on WikiEvents after fine-tuning with human triggers (43.3% F1 pipelined versus 28.4% F1 e2e). We hypothesize that since WikiEvents contains many events per document with its event arguments localized around the trigger (see Table 1), dedicating a system’s component towards trigger extraction, which helps disambiguate events and anchor argument extraction, is particularly beneficial. This is explored in more detail in Section 7.2.
Result #3 – The presence of triggers, even random ones, is needed for prompt-based approaches.
Across a variety of GPT models and multiple datasets, we observe that in-context prompting with triggers (from any trigger source) outperforms no-trigger variations. This is especially drastic for WikiEvents, where we see an over 10 point F1 gain after incorporating triggers in the in-context examples for GPT-4o-mini as shown in Table 2. Interestingly, even adding random spans as triggers boosts performance by  10 points. However, upon manual inspection, we observe that GPT-4o-mini identified “trigger-like" phrases in spite of the nonsense triggers in the in-context examples. This follows findings in recent work Min et al. (2022) and suggests that learning the concept of a trigger, even without proper examples, meaningfully supports extraction.
@@ 7.2 Observations on the Importance of triggers  for Document-Level Event Extraction
We notice in CMNEE and MUC that systems trained without triggers perform comparably to systems trained with triggers (Figure 4(a)). The key commonality between the two datasets is their relatively few number of events per document (compared to WikiEvents). We hypothesize that event disambiguation is therefore easier in MUC and CMNEE; therefore the use of triggers (to anchor argument extraction) for these two datasets is less useful. One exception to this, however, is applying TANL to CMNEE (50.0% F1 in a pipelined setting fine-tuning human triggers versus 40.0% F1 when fine-tuning without any triggers). We believe this may be due to localized events (i.e. we find CMNEE’s arguments also tend to be relatively close together (see Table 1)).
Result #5 – Triggers are needed for corpora that have many localized events to extract in each document.
This result is almost a negation of the previous claim, except for the additional condition of localized events. Aside from the typical number of events per document, WikiEvents also differs from MUC and CMNEE in how closely grouped together event arguments are (see Table 1). We unfortunately do not have a dataset with documents with many events with dispersed arguments to explore the many event per document setting more generally. In Figure 4(b) and Table 2, we see that e2e and pipelined systems for WikiEvents fine-tuned with triggers perform better than their fine-tuned counterparts without triggers. Our conclusion regarding the number of events per document is similar to Result #2: many events require careful event disambiguation and identifying triggers serves this purpose well. Regarding event localization, the trigger’s position is a strong clue to inform systems where to look for associated arguments. In fact, we see that pipelined TANL and TANL + GenIE, the only systems that have access to the exact location of triggers within a document prior to argument extraction, are the best models for WikiEvents (see Table2).
@@
We see in Figure 5(a) that trigger quality does not significantly impact overall performance for fine-tuning methods. This suggests that models can learn textual cues beyond triggers during extraction. An exception to this is the pipelined TANL system applied to WikiEvents, though we believe this relates to the discussion in Section 7.2: in many documents with localized event arguments, learning quality triggers helps disambiguate events during trigger extraction and, in this case, human triggers have proved to be the most useful (relative to other trigger sources) textual anchors for argument extraction.
Result #7 – For pipeline architectures, robustness with respect to trigger quality can be achieved by adding additional event information (e.g. templates, instructions).
We see in Table 2 that e2e variants of systems generally exhibit more consistent performance when varying the fine-tuning examples’ trigger annotation sources. This is especially apparent with the random trigger experiments: in particular TANL is significantly more performant in an e2e setup than in a pipelined setup (40.2% versus 10.9%, 26.9% versus 2.9%, 34.0% versus 17.5% F1 for MUC, WikiEvents, and CMNEE, respectively). However, this difference is remedied by providing a natural language template as input for argument extraction, as in the case of TANL + GenIE: the addition results in +21.3, +11.1, and +31.2 F1 % points for MUC, WikiEvents, and CMNEE, respectively. We observe similar patterns with Degree, which in addition to a template, also provides event type information such as likely triggers and a characterization of typical events: the system does not exhibit any significant swings in overall performance when comparing e2e and pipelined setups or across trigger annotation sources. This suggests that additional relevant inputs such as natural language templates that relate argument types and descriptions of events benefit fine-tuning methods to the extent of possibly overcoming systematic flaws such as conditioning extraction on noise. Overall, we see the trend outlined in Figure 3 where adding templates, then event specific instructions gradually allows system to perform more robustly when training on random triggers.

### v2

@@ 5 Models
We study the performance of four state-of-the-art sequence-to-sequence style systems for document-level event extraction: TANL Paolini et al. (2021), GTT Du et al. (2021), Degree Hsu et al. (2022), and GenIE Li et al. (2021) and compare them against prompt-based baselines using GPT-4o and GPT-4o-mini.
@@ 5.1 Models Overview
We prompt GPT-4o and GPT-4o-mini to extract event templates in a JSON format giving it six in-context examples. We extract all events for all event types using a single query. We use event descriptions and templates as instructions. An example of the prompt can be found in Appendix B.
@@
Paolini et al. (2021) is a pipeline-based model that takes only the document as input. As shown in Figure 2, TANL fine-tunes a T5-large transformer Raffel et al. (2020) to complete event extraction in two phases: 1) event detection, and 2) argument extraction. First, the document is passed through T5 to tag/delimit a trigger for each detected event (the entire document is produced as output with event type tags associated with recognized triggers). Then, each identified {trigger, event type} pair (tagged/delimited in the input document) is passed through T5 once to identify all event arguments and assign them a role type.
@@
Du et al. (2021) is an end-to-end model that takes a document and an ordered list of the possible event types and generates filled templates for all event types in a single pass through BERT Devlin et al. (2019). Templates are linearized to sequences with special argument and event delimiters.
@@
Hsu et al. (2022) is an end-to-end model that takes a document, a natural language template for one event type E and extra information about events of type E (i.e. type definitions, example triggers); it generates filled natural language templates (that include identified triggers) for all detected events for type E in a single pass. The model is called once for each event type of interest for each document. For this reason, training is augmented with multiple negative examples (i.e. document-(empty)template pairs for event types that do not appear in the text).
@@
Li et al. (2021) was originally proposed for argument extraction. Specifically, a document is provided with the trigger delimited by special tokens and a natural language template (similar to that of Degree) corresponding to the event type to extract arguments for. The model output is the natural language template filled with arguments. We use GenIE for the full event extraction task by using TANL as the trigger extractor and GenIE as the argument extractor.
5.2 Model Type: Pipeline vs End-to-End (e2e), No Trigger
We explore both pipeline and end-to-end model architectures and modify TANL and Degree to run in both settings. As shown in Figure 2, a pipeline system consists of a trigger extraction and an argument extraction component. The trigger extraction phase identifies all event triggers and their corresponding event types. The argument extraction phase then operates on each {trigger, event type} pair separately. This design allows argument extraction to anchor in an informative part of the text, i.e. in the vicinity of a known trigger. End-to-End systems, on the other hand, jointly extract triggers, event types, and arguments, allowing the two extraction tasks to inform one another, thus alleviating cascading errors that can occur in a pipelined setting.
Pipeline
We modify Degree to fit the pipeline architecture. We train distinct models for trigger and argument extraction. Specifically, the trigger extraction input and output templates only contain slots for a trigger and event type; the argument extraction input templates have an additional slot for a trigger (identified during trigger extraction) and the output templates only remove the slot for the trigger, retaining only slots for each argument type. The model is called once to fill natural language templates for all identified triggers (trigger extraction) and once for each {identified trigger, event type} pair to identify arguments (argument extraction). We did not use GTT for pipelined experiments. By default, TANL, TANL + GenIE are pipelined models.
No Triggers
We modify all models to run without the use of triggers. All no trigger experiments were conducted in an e2e manner. For TANL, the raw text is provided and the model tags arguments with an event type and number and the argument role (e.g. “perpetrating individual for attack 1”). A single span can have multiple such tags if it is a shared entity mention for multiple events. The spans associated with each unique {event type, number} pair are processed into fillers for single events. This scheme only requires one pass of T5 for full event extraction. For Degree, we only remove the slot for the trigger in input/output templates. By default, GTT is an e2e system that does not use triggers.
End-to-End (with triggers)
For TANL, the event type and number and the argument role tag is supplemented with a tag for the trigger (e.g. “perpetrating individual for attack 1 | trigger = killed”). There is still only one pass for full event extraction. For GTT, a slot is added at the start of the output event templates for a trigger. By default, Degree is an e2e system using triggers.
6 Experimental Setup
In our study of the utility of triggers on document-level event extraction, we tested the performance of fine-tuned (i.e. pipeline and end-to-end) and prompt-based approaches on the three datasets described in Section 4. We consider the prompt-based approaches as a baseline and compare them to fine-tuning the four state-of-the-art models, described in Section 5; each model was tested in a pipeline and an end-to-end manner as applicable. We train on 1 $\times$ A100 and 4 $\times$ 1080Ti using the default learning rate suggested for each model using the first 10% of the steps as warmup. The reported score on the test data is obtained using the system from the best performing epoch using the validation set.
Due to the varying sizes of the datasets, with some very small, we perform 5-cross validation on the test data by sampling 80% of the data for each evaluation run. Each model is trained once and evaluated once on each of the 5 validation splits. We then report average and standard deviation scores across the 5 splits. Our evaluation focuses on Micro-F1 — combined F1 score on both event type detection and argument role extraction. Since the datasets are multi-event, we have to align the predicted and gold templates. We follow the matching algorithm of Das et al. (2022) to align and score the predicted templates [^6] We modify their proposed scorer by improving its compute time (using the Hungarian Matching algorithm Kuhn (1955)) and by removing span-based scoring as not all models predict spans..
7 Results and Discussion
In this section, we discuss our experimental results shown in Table 2 (follows Conclusions). We first analyze some high level performance trends independent of trigger usage before commenting on how triggers should be used depending on various system and corpora specifications.
7.1 General Observations for Document-Level Event Extraction
As shown in Table 2, the best prompt-based systems achieve 43.8% (with human triggers), 38.3% (keyword), and 66.7% (human) F1 while the best fine-tuning methods achieve 59.7% ($\textsc{{DEGREE}}_{e2e}$ no-trigger), 47.1% ($\textsc{{TANL}}_{pipeline}$ human), and 78.9% F1 ($\textsc{{DEGREE}}_{e2e}$ no trigger) on MUC, WikiEvents and CMNEE, respectively.
@@ 7.1 General Observations for Document-Level Event Extraction
Figure 3: E2E vs. Pipeline. F1-score of each e2e experiment vs. its pipeline version. If the data point is above $y=x$ (dashed line), then the e2e architecture outperformed the pipeline architecture for that given (model, trigger type, dataset) setting. (Datasets cannot be distinguished in the figure.)
As shown in Figure 3, e2e models trained with triggers perform roughly on par or better than their pipelined counterparts with five explainable exceptions. All exceptions involve the TANL system applied to WikiEvents and CMNEE. We hypothesize that since event arguments are localized around the trigger for both datasets (see Table 1), dedicating a separate pipeline component towards trigger extraction, which helps disambiguate events and anchor argument extraction, is particularly beneficial. (This is explored in more detail in Section 7.2 where the propensity of Degree to ignore triggers is also discussed.) This is important because we can generally build e2e systems that are faster to train and quicker during inference than pipelined systems without sacrificing performance.
Result #3 – The presence of triggers, even random ones, is generally needed for prompt-based approaches.
Across both GPT models and multiple datasets, we observe that in-context prompting with triggers (from any trigger source) regularly outperforms the associated no-trigger variations. This is especially drastic for GPT-4o on WikiEvents, where we see a 5 point F1 gain after incorporating keyword triggers in the in-context examples as shown in Table 2. Interestingly, even adding random spans as triggers boosts performance by  2 points for GPT-4o. Upon manual inspection we observe that GPT-4o actually identifies legitimate trigger-like phrases in spite of the nonsense triggers in the in-context examples. This follows findings in recent work Min et al. (2022) and suggests that learning the concept of a trigger, even without proper examples, meaningfully supports extraction.
@@ 7.2 Observations on the Importance of triggers  for Document-Level Event Extraction
Figure 4: Triggers vs. no-Triggers. F1-score of each fine-tuning experiment with triggers (annotated by humans, LLMs, or the keyword heuristic) against its e2e version without triggers. A data point above $y=x$ (dashed line) implies that fine-tuning with triggers performed better for that (model, architecture, dataset).
@@
We notice for CMNEE and (especially) MUC that systems trained without triggers perform comparably to systems trained with triggers (Figure 4(a)) (i.e. most points fall below the $x=y$ diagonal). The key commonality between the two datasets is their relatively low number of events per document (compared to WikiEvents). We hypothesize that event disambiguation is therefore easier in MUC and CMNEE; therefore the use of triggers (to anchor argument extraction) for these two datasets is less useful. Exceptions to this occur when applying TANL to CMNEE where fine-tuning in the pipelined setting performs better with triggers (e.g., 54.3%, 59.6% and 64.0% F1) than without (43.8% F1). We believe this may be due to localized events for CMNEE as noted in Result #2 and Result #5.
Result #5 – Higher quality triggers are preferred for corpora that have many localized events to extract in each document.
This result is almost a negation of the previous claim, except for the additional conditions of localized events and higher quality triggers. Aside from the typical number of events per document, WikiEvents also differs from MUC and CMNEE in how closely grouped together event arguments are (see Table 1). We unfortunately do not have a dataset with documents with many events with dispersed arguments to explore the many event per document setting more generally. In Figure 4(b) and Table 2, we see that e2e and pipelined systems for WikiEvents fine-tuned with human triggers perform better than their fine-tuned counterparts without triggers. Our conclusion regarding the number of events per document is similar to Result #2: many events require careful event disambiguation and identifying relevant triggers serves this purpose well. Regarding event localization, the trigger’s position is a strong clue to inform systems where to look for associated arguments. In fact, we see that pipelined TANL and TANL + GenIE, the only systems that have access to the exact location of triggers within a document prior to argument extraction, are the best models for WikiEvents (see Table2).
@@
We see in Figure 5(a) that trigger quality does not significantly impact overall performance for fine-tuning methods. This suggests that models can learn textual cues beyond triggers during extraction. An exception to this is the pipelined TANL system applied to WikiEvents, though we believe this relates to the discussion for Result #5.
Result #7 – Robustness with respect to degraded trigger quality can be achieved by giving the event extraction system access to additional event information (e.g. templates, instructions).
Evidence for this can be seen by examining the performance of Degree, which does not exhibit any significant swings in overall performance when comparing across trigger annotation sources. For MUC and CMNEE, Degree actually performs best (and outperforms all other systems) in the no triggers setting (59.7% and 78.9% F1, respectively). We believe that this robustness arises from the background information it requires as part of its input, including a natural language statement of the template for the target event type, a definition of the event type, a short list of likely triggers, short definitions of expected event arguments. As additional evidence for this claim, one can observe the random triggers results for TANLpipeline in Table 2, which are significantly worse than its e2e variation (11.6% versus 47.3%, 5.9% versus 34.3%, 22.3% versus 39.4% F1 for MUC, WikiEvents, and CMNEE, respectively). This difference, however, is remedied by providing a natural language template as input for argument extraction, as in the case of TANL + GenIE: the addition results in +20.6, +6.1, and +26.4 F1 % points for MUC, WikiEvents, and CMNEE, respectively. These results suggest that additional relevant inputs such as natural language templates that relate argument types and descriptions of events benefit fine-tuning methods to the extent of possibly overcoming systematic flaws such as conditioning extraction on noise. Overall, we see the trend outlined in Figure 5(b): TANL exhibits very poor performance, adding event template information (TANL + GenIE) improves system performance, and the inclusion of event-specific instructions (Degree) allows further robustness when training on random triggers.